# In-Run Data Shapley for Adam Optimizer

### v2

@@ 1 Introduction
Modern machine learning models rely heavily on large-scale training data, yet model performance and behavior are typically governed by the quality and influence of individual samples. The lack of reliable data attribution mechanisms can lead to degraded performance and computational inefficiency. Without principled mechanisms to isolate and evaluate sample influence, trained models may amplify harmful biases or remain vulnerable to data poisoning attacks (Koh and Liang, 2017; Ghorbani and Zou, 2019; Ren et al., 2025a). Furthermore, the presence of problematic data escalates the computational cost of modern training pipelines by wasting resources on uninformative or harmful examples (Paul et al., 2021; Ren et al., 2025b; Hu et al., 2025).
To fairly attribute data contributions, the Shapley value has emerged as the gold standard. Originating from cooperative game theory (Shapley, 1953), the Shapley value provides a unique, unbiased distribution of the total model performance among individual data points based on their marginal contributions, emerging as a widely adopted framework (Ghorbani and Zou, 2019; Jia et al., 2019a). While it has desirable properties, such as fairness, efficiency, and additivity (Jia et al., 2019b), the classic calculation of Shapley values is computationally prohibitive for deep learning. Formally, it requires retraining the model exponentially many times using different subsets of data to evaluate the marginal contribution of each source. This complexity has limited the exact application of Shapley values to only the smallest datasets and models, necessitating approximation methods for practical use.
To overcome this scalability bottleneck, recent work has introduced In-Run Data Shapley (Wang et al., 2025). This method fundamentally shifts the attribution paradigm by estimating Shapley values dynamically during a single training process, rather than through retrospective retraining. However, despite its efficiency, the existing In-Run Data Shapley is specifically tailored for Stochastic Gradient Descent (SGD) (Robbins and Monro, 1951), and not directly extendable to other popular optimizers like Adam (Kingma and Ba, 2014). However, in practice, modern deep learning models are almost exclusively trained using adaptive optimizers, most notably Adam. This discrepancy motivates us to ask the following questions:
@@ 3 Background
Adapting the concept from game theory (Shapley, 1953), we treat the data points as players in a game and the model performance as the total payout. The Data Shapley value, denoted as $\phi_{z}(U)$, provides a unique and fair way to distribute the total utility among the $N$ data points.
@@
Lemma 3.1 (Linearity of the Shapley value (Shapley, 1953)).
@@
Adam Optimizer and State Coupling. Adam (Kingma and Ba, 2014) is a widely used adaptive optimizer that maintains exponential moving averages of both first-order ($m_{t}$) and second-order ($v_{t}$) gradient moments. At each step $t$, the update is given by:
@@ 4 In-Run Data Shapley for Adam Optimizer
In this section, we first demonstrate that In-Run Data Shapley (Wang et al., 2025) is inherently dependent on the choice of optimization algorithm. Motivated by this discrepancy, we then propose an optimizer-aware extension tailored specifically for Adam-based training.
@@ 4.1 Optimizer Dependence of Data Value
To ensure our findings are robust and reflect true marginal contributions rather than artifacts of a specific approximation method, we relied on Truncated Monte Carlo (TMC) (Ghorbani and Zou, 2019). TMC estimates the Shapley value by retraining the model on random subsets of data, accumulating the marginal validation loss changes induced by individual samples along the optimization trajectory. Although computationally expensive, this retraining-based approach serves as a model-agnostic “ground truth” baseline, allowing us to rigorously compare how the optimization path itself alters data value.
@@ 4.2 In-Run Data Shapley for Adam
Standard In-Run Data Shapley (Wang et al., 2025) assesses the value of a training set by decomposing its total utility into contributions accumulated across iterations. Instead of retraining, it evaluates a local utility function at each step. Formally, given a validation data point $z^{\text{(val)}}$ and a sequence of model checkpoints $\{w_{t}\}_{t=0}^{T}$, the local utility function at a single iteration $t$ can be defined as:
@@
Extending this to Adam is non-trivial due to the stateful and non-linear nature of its update rule, as noted by (Wang et al., 2025). By applying a first-order Taylor expansion to the local utility function under Adam dynamics, we can derive a tractable, closed-form estimator that restores the additive property required for Shapley valuation:
@@
In-Run Data Shapley decomposes the overall utility into step-wise contributions, a formulation that also benefits the Adam optimizer, as illustrated in Theorem 4.1. However, compared to (Wang et al., 2025), our result differs in that the Data Shapley value of a training point accumulates dot products between its update directions and the gradients of the validation data, rather than gradient–gradient dot products as in Wang et al. (2025). This distinction precludes the use of the efficient gradient dot-product computation techniques employed in prior work, a challenge we address in the subsequent section.
@@ 4.3 Efficient Computation via Ghost Dot-Product
Standard In-Run Data Shapley (Wang et al., 2025) achieves efficient pairwise gradient dot products between each training data and the validation point by Ghost Dot-Product. Specifically, directly computing the inner product $\nabla\ell(w_{t},z_{i})^{\top}\nabla\ell(w_{t},z^{(\text{val})})$ would require explicitly instantiating per-sample gradients for the entire parameter space, incurring a prohibitive memory cost of $O(B\times P)$, where $B$ is the batch size and $P$ is the number of model parameters.
@@
$\nabla\ell(w_{t},z_{i})^{\top}\nabla\ell(w_{t},z^{(\text{val})})$ $=\sum_{l}\left\langle\delta_{i}^{(l)}(a_{i}^{(l)})^{\top},\delta_{\text{val}}^{(l)}(a_{\text{val}}^{(l)})^{\top}\right\rangle_{F}$
$=\sum_{l}\underbrace{(\delta_{i}^{(l)\top}\delta_{\text{val}}^{(l)})}_{\text{Error correlation}}\cdot\underbrace{(a_{i}^{(l)\top}a_{\text{val}}^{(l)})}_{\text{Activation correlation}}.$
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v3

@@ 1 Introduction
Modern machine learning models rely heavily on large-scale training data, yet model performance and behavior are typically governed by the quality and influence of individual samples. The lack of reliable data attribution mechanisms can lead to degraded performance and computational inefficiency. Without principled mechanisms to isolate and evaluate sample influence, trained models may amplify harmful biases or remain vulnerable to data poisoning attacks Koh and Liang (2017); Ghorbani and Zou (2019); Ren et al. (2025a). Furthermore, the presence of problematic data escalates the computational cost of modern training pipelines by wasting resources on uninformative or harmful examples Paul et al. (2021); Ren et al. (2025b); Hu et al. (2025).
To fairly attribute data contributions, the Shapley value has emerged as the gold standard. Originating from cooperative game theory Shapley (1953), the Shapley value provides a unique, unbiased distribution of the total model performance among individual data points based on their marginal contributions, emerging as a widely adopted framework Ghorbani and Zou (2019); Jia et al. (2019a). While it has desirable properties, such as fairness, efficiency, and additivity (Jia et al., 2019b), the classic calculation of Shapley values is computationally prohibitive for deep learning. Formally, it requires retraining the model exponentially many times using different subsets of data to evaluate the marginal contribution of each source. This complexity has limited the exact application of Shapley values to only the smallest datasets and models, necessitating approximation methods for practical use.
To overcome this scalability bottleneck, recent work has introduced In-Run Data Shapley Wang et al. (2025). This method fundamentally shifts the attribution paradigm by estimating Shapley values dynamically during a single training process, rather than through retrospective retraining. However, despite its efficiency, the existing In-Run Data Shapley is specifically tailored for Stochastic Gradient Descent (SGD) Robbins and Monro (1951), and not directly extendable to other popular optimizers like Adam Kingma and Ba (2014). However, in practice, modern deep learning models are almost exclusively trained using adaptive optimizers, most notably Adam. This discrepancy motivates us to ask the following questions:
@@ 3 Background
Adapting the concept from game theory Shapley (1953), we treat the data points as players in a game and the model performance as the total payout. The Data Shapley value, denoted as $\phi_{z}(U)$, provides a unique and fair way to distribute the total utility among the $N$ data points.
@@
Lemma 3.1 (Linearity of the Shapley value Shapley (1953)).
@@
Adam Optimizer and State Coupling. Adam Kingma and Ba (2014) is a widely used adaptive optimizer that maintains exponential moving averages of both first-order ($m_{t}$) and second-order ($v_{t}$) gradient moments. At each step $t$, the update is given by:
@@ 4 In-Run Data Shapley for Adam Optimizer
In this section, we first demonstrate that In-Run Data Shapley Wang et al. (2025) is inherently dependent on the choice of optimization algorithm. Motivated by this discrepancy, we then propose an optimizer-aware extension tailored specifically for Adam-based training.
@@ 4.1 Optimizer Dependence of Data Value
To ensure our findings are robust and reflect true marginal contributions rather than artifacts of a specific approximation method, we relied on Truncated Monte Carlo (TMC) Ghorbani and Zou (2019). TMC estimates the Shapley value by retraining the model on random subsets of data, accumulating the marginal validation loss changes induced by individual samples along the optimization trajectory. Although computationally expensive, this retraining-based approach serves as a model-agnostic “ground truth” baseline, allowing us to rigorously compare how the optimization path itself alters data value.
@@ 4.2 In-Run Data Shapley for Adam
Standard In-Run Data Shapley Wang et al. (2025) assesses the value of a training set by decomposing its total utility into contributions accumulated across iterations. Instead of retraining, it evaluates a local utility function at each step. Formally, given a validation data point $z^{\text{(val)}}$ and a sequence of model checkpoints $\{w_{t}\}_{t=0}^{T}$, the local utility function at a single iteration $t$ can be defined as:
@@
Extending this to Adam is non-trivial due to the stateful and non-linear nature of its update rule, as noted by Wang et al. (2025). By applying a first-order Taylor expansion to the local utility function under Adam dynamics, we can derive a tractable, closed-form estimator that restores the additive property required for Shapley valuation:
@@
In-Run Data Shapley decomposes the overall utility into step-wise contributions, a formulation that also benefits the Adam optimizer, as illustrated in Theorem 4.1. However, compared to Wang et al. (2025), our result differs in that the Data Shapley value of a training point accumulates dot products between its update directions and the gradients of the validation data, rather than gradient–gradient dot products as in Wang et al. (2025). This distinction precludes the use of the efficient gradient dot-product computation techniques employed in prior work, a challenge we address in the subsequent section.
@@ 4.3 Efficient Computation via Ghost Dot-Product
Standard In-Run Data Shapley Wang et al. (2025) achieves efficient pairwise gradient dot products between each training data and the validation point by Ghost Dot-Product. Specifically, directly computing the inner product $\nabla\ell(w_{t},z_{i})^{\top}\nabla\ell(w_{t},z^{(\text{val})})$ would require explicitly instantiating per-sample gradients for the entire parameter space, incurring a prohibitive memory cost of $O(B\times P)$, where $B$ is the batch size and $P$ is the number of model parameters.
@@
$\nabla\ell(w_{t},z_{i})^{\top}\nabla\ell(w_{t},z^{(\text{val})})$ $=\sum_{l}\left\langle\delta_{i}^{(l)}(a_{i}^{(l)})^{\top},\delta_{\text{val}}^{(l)}(a_{\text{val}}^{(l)})^{\top}\right\rangle_{F}=\sum_{l}\underbrace{(\delta_{i}^{(l)\top}\delta_{\text{val}}^{(l)})}_{\text{Error correlation}}\cdot\underbrace{(a_{i}^{(l)\top}a_{\text{val}}^{(l)})}_{\text{Activation correlation}}.$